How many liters of a 10% solution needs to be added to 30 liters of a 60% solution to get a 40% solution?
Answer by Earlsdon(6294)

Let x = the number of liters of 10% solution to be added to the 30 liters of the 60% solution. Change the percentages to their decimal equivalents.
0.1x+0.6(30) = 0.4(30+x)
0.1x+18 = 12+0.4x Subtract 12 from both sides.
0.1x +6 = 0.4x Subtract 0.1x from both sides.
6 = 0.3x Finally, divide both sides by 0.3
20 = x
You will need to add 20 liters of 10% solution to the 30 liters of 60% solution to obtain 50 liters of 40% solution.